∫∫∫cos(x^2+y^2+z^2)^(3/2) dxdydz
1
Expert's answer
2020-03-03T11:44:06-0500

"\iiint{\cos(x^2+y^2+z^2)^{3/2} dxdydz}"


Convert "x,y,z" to spherical coordinates "r, \varphi, \theta."

"x=r\cos(\varphi)\sin(\theta) \\
y=r\sin(\varphi)\sin(\theta) \\
z=r\cos(\theta)"


The absolute value of Jacobian is "r^2\sin(\theta)." Then the original integral is



"\iiint {\cos((r^2)^{3/2})\cdot r\sin(\theta)dr d\varphi d\theta} = \iiint{r^2\cos(r^3)\sin(\theta)dr d\varphi d\theta } \\

= \int{r^2\cos(r^3)dr}\int{d\varphi}\int{\sin(\theta)d\theta} = \frac{1}{3}\int{\cos(r^3)d(r^3)}\int{d\varphi}\int{\sin(\theta)d\theta} \\

= -\frac{1}{3}\varphi\sin(r^3)\cos{(\theta)}"

We have omitted the constants for convenience.

Return to "x,y,z."

"r=\sqrt{x^2+y^2+z^2} \\
\varphi = \arctan{\left( \frac{y}{x}\right)} \\
\theta = \arccos{\left(\frac{z}{\sqrt{x^2+y^2+z^2}}\right)}"


Therefore, the original integral is


"\iiint{\cos(x^2+y^2+z^2)^{3/2} dxdydz} \\
= -\frac{1}{3} \arctan{\left( \frac{y}{x}\right)}\sin((x^2+y^2+z^2)^{3/2})\cos{\left[ \arccos{\left(\frac{z}{\sqrt{x^2+y^2+z^2}}\right)} \right]} \\

= -\frac{1}{3} \arctan{\left( \frac{y}{x}\right)}\sin((x^2+y^2+z^2)^{3/2})\frac{z}{\sqrt{x^2+y^2+z^2}}"